OpenStudy (anonymous):

Help I don't get this stuff 2. The table below represents a linear function f(x) and the equation represents a function g(x). X F(x) 0 28 1 30 2 32 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points) Part A: Find the Slope of f(x): Find the Slope of g(x): Write a sentence comparing the two slopes: Part B: What is the y-intercept of f(x): What is the y-intercept of g(x): Which one is greater?

OpenStudy (anonymous):

Please help im so confused ill give a medal

OpenStudy (aum):

g(x) = ? g(x) is missing in the problem.

OpenStudy (anonymous):

Im still confused

OpenStudy (aum):

You have not copied the problem fully. The problem states: "the equation represents a function g(x)" But I don't see an equation for g(x). You have not copied the part g(x) = ....

OpenStudy (anonymous):

g(x) = 10x - 25

OpenStudy (aum):

x f(x) 0 28 1 30 2 32 Slope = \(\Large \frac{y_2-y_1}{x_2-x_1} = \frac{30-28}{1-0} = 2\). Slope of f(x) = 2 g(x) = 10x - 25. Compare it to y = mx + b where m is the slope. m = 10. Slope of g(x) = 10 Part A: Find the Slope of f(x): 2 Find the Slope of g(x): 10 Write a sentence comparing the two slopes: g(x) has a greater slope and therefore steeper than f(x).

OpenStudy (anonymous):

Thank you what would B be at the bottom part

OpenStudy (aum):

y-intercept is the y-value when x = 0. For f(x), the table shows that when x = 0, y or f(x) = 28. Therefore, the y-intercept of f(x) is 28. g(x) = 10x - 25 y-intercept is when x = 0. put x = 0 g(x) = 0 - 25 = -25 Therefore, the y-intercept of g(x) is -25. 28 > -25. Therefore, f(x) has a greater y-intercept than g(x).